OpenStudy (anonymous):

Water coming out from a fountain is modeled by the function f(x) = -x2 + 5x + 4 where f(x) represents the height, in feet, of the water from the fountain at different times x, in seconds. What does the average rate of change of f(x) from x = 3 to x = 5 represent? The water falls down with an average speed of 3 feet per second from 3 seconds to 5 seconds. The water falls down with an average speed of 5 feet per second from 3 seconds to 5 seconds. The water travels an average distance of 3 feet from 3 seconds to 5 seconds. The water travels an average distance of 5 feet from 3 seconds

OpenStudy (freckles):

average rate of change is given by the change of y over the change of x. \[\frac{ f(5)-f(3)}{5-3} \]

OpenStudy (freckles):

you are dividing a change of height by a change in time

OpenStudy (anonymous):

So x = 2?

OpenStudy (dumbcow):

also "rate of change" = speed

OpenStudy (anonymous):

So, I know its either A or B, right?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

So its either 5 or 3 feet per 3-5 seconds?

OpenStudy (freckles):

just used the formula I gave above to figure out between A and B

OpenStudy (anonymous):

I dont understand the formula you gave me.

OpenStudy (anonymous):

f(5) - f(3) = f(2) ----- 2 ??????????????

OpenStudy (freckles):

...how did you get f(5)-f(3)=f(2)?

OpenStudy (freckles):

f(5)=-5^2+5(5)+4 and f(3)=-3^2+5(3)+4

OpenStudy (anonymous):

So f(5) = 3154 And f(3) = 110

OpenStudy (freckles):

-5^2+5(5)+4 ------ first we know 5^2=5(5)=25 so -5^2+5(5)+4 can be written as -25+25+4

OpenStudy (freckles):

-25+25 is 0 though so you have -5^2+5(5)+4 can be written as 0+4 which is just 4

OpenStudy (freckles):

can you try doing f(3)

OpenStudy (anonymous):

f(3) = 25?

OpenStudy (freckles):

are you guessing?

OpenStudy (anonymous):

No

OpenStudy (freckles):

f(x)=-x^2+5x+4 just plug in 3 where x is

OpenStudy (anonymous):

OH there is a negative, my bad

OpenStudy (freckles):

f(3)=-3^2+5(3)+4 3^2 means 3(3) which is 9 5(3) is 15 f(3)=-9+15+4

OpenStudy (anonymous):

so 10?

OpenStudy (freckles):

\[\frac{f(5)-f(3)}{5-3}=\frac{4-10}{5-3}\]

OpenStudy (anonymous):

so its 3

OpenStudy (anonymous):

So the answer is a?

OpenStudy (freckles):

well -3 where the units is distance/time which is a speed unit which just means going down at a speed of 3ft/sec

OpenStudy (freckles):

and yes a sounds great

OpenStudy (anonymous):

Thx for the help! Ima submit my test now, lets see :D

OpenStudy (anonymous):

Oh no! I got a %62 on my test :(. 5 questions wrong out of 13. But the one you helped me with was right, ty so much!